Trial 1 Volume of acid 30 ml Volume NaOH ADDED TO HALF NEUTRALIZE ACID 6.5ML Ph...

Trial 1

Volume of acid

30 ml

Volume NaOH ADDED TO HALF NEUTRALIZE ACID

6.5ML

Ph of Half neutralized solution

4.34

Ka

?

PKa of acetic acid

?
Answer #1

We know that, pH corresponding to half equivalence point = pKa

In this case, pH corresponding to half equivalence point = 4.34

Therefore, pKa of acetic acid = 4.34

We have relation, pKa = - log 10 Ka

\therefore Ka = 10 - pKa  = 10 -4.34 = 4.57 \times 10 -05

ANSWER : pKa of acetic acid = 4.34 & Ka of acetic acid = 4.57 \times 10 -05